# Networked Multiagent Reinforcement Learning for Peer-to-Peer Energy Trading

### v2

@@ Abstract
Utilizing distributed renewable and energy storage resources in local distribution networks via peer-to-peer (P2P) energy trading has long been touted as a solution to improve energy systems’ resilience and sustainability. Consumers and prosumers (those who have energy generation resources), however, do not have the expertise to engage in repeated P2P trading, and the zero-marginal costs of renewables present challenges in determining fair market prices. To address these issues, we propose multi-agent reinforcement learning (MARL) frameworks to help automate consumers’ bidding and management of their solar PV and energy storage resources, under a specific P2P clearing mechanism that utilizes the so-called supply-demand ratio. In addition, we show how the MARL frameworks can integrate physical network constraints to realize voltage control, hence ensuring physical feasibility of the P2P energy trading and paving way for real-world implementations.
Introduction
@@
The rest of the paper is organized as follows. Section II reviews the literature on P2P energy trading and emphasizes the contribution of our work. Section III introduces the market clearing mechanism based on the supply-demand ratio, first proposed in [5], which can avoid zero-clearing prices when the supply resources have zero marginal costs. Section IV provides the detailed consensus-MARL formulation and algorithm, as well as establishing convergence results. The specific setups and data of our simulation are described in Section V, along with numerical results. Finally, Section VI summarizes our work and points out several future research directions.
Literature Review
@@
The literature on decentralized control of energy transactions in a distribution network is extensive and diverse. Broadly speaking, there are two predominant approaches: the distributed optimization approach and P2P bilateral trading. A key method in the former approach is the alternating direction multiplier method (ADMM), which solves centralized dispatch problems by iteratively updating price signals for consumer/prosumer optimization (such as [6, 7, 8]). Despite ADMM’s theoretical convergence under convexity, practical challenges arise: (i) As an algorithmic approach rather than a market design, ADMM is subject to the same difficulties faced by a uniform-pricing-based market of all zero-marginal-cost resources. Specifically, it is prone to ‘bang-bang’ pricing: market clearing prices plummet to zero in oversupply or spike to a price ceiling in undersupply. Addressing this with additional market mechanisms, such as adding ancillary service markets, akin to wholesale markets, could overly complicate distribution-level markets for energy trading. (ii) Most distributed algorithms, including ADMM, are static, lacking real-time adaptability. While online ADMM variants exist [9], ensuring network constraints within each iteration remains an open question. (iii) Assuming consumers or prosumers can solve complex optimization problems is often unrealistic due to potential expertise or resource limitations.
@@
These limitations form the basis of our interest in exploring P2P bilateral trading. Consequently, our literature review will focus on this alternative approach. There have been multiple review papers on the large and ever-growing literature on P2P energy trading [10, 11, 12, 13]. Within the literature, we focus on the works that explicitly consider distribution network constraints. [14] proposes a continuous double auction framework within which physical network constraints are maintained through a sensitivity-based approach. The agents in the auction are the so-called zero-intelligence-plus (ZIP) traders, who employ simple adaptive mechanisms without any learning of the repeated multi-agent interactions. More sophisticated than ZIP, a fully decentralized MARL is adopted in [15] to realize decentralized voltage control and is extended in [2] to form a P2P energy trading market with physical constraints. These two papers implement a completely decentralized approach in the sense that each agent ignores the multi-agent interaction as well and just learns their single-agent policy to maximize their own payoff.
@@
Another MARL approach that has been explored is the centralized training and decentralized execution (CTDE) framework, of which the multi-agent deep deterministic policy gradient (MADDPG) method [4] is a prominent example. It is applied in [16] to solve the autonomous voltage control problem In [17], a modified MADDPG algorithm is proposed for a double-sided auction market to facilitate P2P energy trading, although physical constraints are not considered in their model. While the CTDE framework can lead to more efficient learning and robust outcomes, its dependence on a centralized entity for collecting and coordinating all agents’ data introduces difficulties for practical implementations and potential vulnerabilities, such as increased risk of single-point failures and higher susceptibility to cyber-attacks
@@
In this work, we present several key contributions: First, we develop a comprehensive framework for a P2P energy trading market while integrating a consensus-based MARL algorithm. This framework facilitates automated control and bidding of DERs while allowing for decentralized learning of voltage constraints. Second, we present the sufficient conditions required for the convergence of the MARL algorithm and show the details of implementing this algorithm using deep neural networks. Third, we conduct a comparative analysis of market outcomes derived from three distinct MARL strategies: a naive decentralization approach, the MADDPG approach, and the proposed consensus-based MARL. Our findings reveal that the consensus-based approach notably outperforms the others in achieving better market outcomes.
Market Clearing Mechanism
@@
$P_{t}:=P(SDR_{t})$
:=$:=\begin{cases}(FIT-UR)\cdot SDR_{t}+UR,&0\leq SDR_{t}\leq 1;\\ FIT,&SDR_{t}>1.\end{cases}$ (2)
In (III), $P_{t}$ denotes the market clearing price in round $t$, which is a piece-wise linear function with respect to the $SDR_{t}$, as illustrated in Figure 1.
@@
While the SDR approach may be criticized for not being based on economic theories, we argue that from our perspectives, the simplicity and transparency of the SDR mechanism outweigh the criticisms, and it is well-suited for a P2P energy market with all-zero marginal cost resources. In our previous work [20], it is identified that a P2P energy market using traditional double-sided auction may not work well due to the special characteristics of such a market. Since uncleared electricity demand bids have to be bought at the rate of $UR$, and uncleared energy supply bids have to be sold at $FIT$, with the $UR$ and $FIT$ known to all market participants. With all zero-marginal cost resources, agents would not know how to price their bids in a marginal-cost-based clearing mechanism, and the public information of the $UR$ and the $FIT$ naturally provides the focal points for sellers to bid at $UR$ and buyers to bid at $FIT$, aiming to maximize their own payoffs. This would lead to a bang-bang type market outcome: when there is overdemand, the market price would be UR; reversely, when there is oversupply, the price would be FIT. We view the SDR approach as a fair and reasonable alternative to the marginal-cost-based approach to avoid such bang-bang outcomes, especially for a new market in which the supply resources are likely much less than demand.
@@
Additionally, the SDR method offers two key benefits: First, it maintains the market clearing price between $UR$ and $FIT$, ensuring that transactions for buyers and sellers are at least as favorable as dealing directly with the utility. Second, it simplifies the bidding process by requiring only quantity bids, avoiding the complexity of submitting bids in price-quantity pairs.
Decentralized Learning in P2P Trading with Consensus Updates
@@
Markov Decision Process for a Prosumer
@@
$e_{i,t+1}=$
$\max\Big{\{}\min\Big{[}e_{i,t}+\eta^{c}_{i}\max(a^{e}_{i,t},0)+\frac{1}{\eta^{d}_{i}}\min(a^{e}_{i,t},0),\overline{e}_{i}\Big{]},0\Big{\}},$ (3)
where $\eta^{c}_{i}$ and $\eta^{d}_{i}$ are the charging and discharging efficiency of agent $i$’s battery, respectively, and $\overline{e}_{i}$ is the battery capacity. The charging efficiency represents the ratio of the amount of energy effectively stored in the system to the energy input during the charging process; while the discharging efficiency is the ratio of the energy delivered by the system during discharge to the energy that was initially stored in it. Their product, $\eta^{c}_{i}\times\eta^{d}_{i}$, yields the so-called round-trip efficiency. The outside “max” operation in (IV-A) is to ensure that the energy level in the battery will not be negative; while the first “min” operation in the bracket is to ensure that the battery storage capacity is not exceeded.
Reward: Each agent $i$’s reward function is affected by both the collective actions of the agents and the system states. Specifically, agent $i$’s reward in time step $t$, denoted by $r_{i,t}$, is a random variable whose conditional expectation has two components as follows:
$\begin{split}\mathbb{E}[r_{i,t}|s_{t},a_{t}]&:=R_{i}(a_{i,t};a_{-i,t},s_{t})\\ &\ =R^{m}_{i}(a^{e}_{i,t};a^{e}_{-i,t},s_{t})+R^{v}(a_{i,t};a_{-i,t},s_{t})/I.\end{split}$ (4)
@@
which is simply the net energy of PV generation minus baseload demand (of real power) and charge to the battery. The market clearing price $P_{t}$ in (5) is defined in (III).
@@
The second term in the reward function, $R^{v}$, is the system voltage violation penalty at a given time $t$. To provide an explicit form of $R^{v}$, we first write out the standard bus-injection model. Assume that the network has $N$ buses, we have that for $\kappa=1,\ldots,N$ (the time index $t$ is omitted below for simplicity):
@@
In (8), $\overline{V}_{\kappa}$ and $\underline{V}_{\kappa}$ represent the upper and lower limit of bus $k$’s voltage magnitude, respectively. Let $f_{clip}$ denote a generic ‘clip’ function in the form of $f_{clip}=\text{clip}[x,\{v_{min},v_{max}\}]$, which is defined as $f_{clip}=x$ when $v_{min}\leq x\leq v_{max}$, $f_{clip}=v_{min}$ if $x<v_{min}$, and $f_{clip}=v_{max}$ if $x>v_{max}$. Based on the clip function definition, if the voltage magnitudes $|V_{\kappa,t}|$ obtained from solving equations (7) are within the voltage limit of $[\underline{V}_{\kappa},\overline{V}_{\kappa}]$ on every bus $\kappa=1,\ldots,N$, then $R^{v}=0$; if there is voltage violation at a certain bus $\kappa$, the term $\max(|V_{\kappa,t}|-\overline{V}_{\kappa},\ \underline{V}_{\kappa}-|V_{\kappa,t}|)$ becomes positive, and $R^{v}$ becomes a negative number with an arbitrary positive number $\lambda$ to amplify the voltage violation. The other positive parameter in (7), $M$, is just to ensure that the reward $R^{v}$ is bounded below, a technical condition needed to ensure the convergence of the consensus MARL algorithm to be introduced in the next section.
If (7) does not have a solution with a given set of agents’ bids, we can simply set $R^{v}=-\lambda\cdot N\cdot M$. We want to emphasize that the voltage violation does not result in actual financial penalties for the agents since the market clearing happens before real-time delivery. It is simply feedback sent to agents for them to train their reinforcement learning policies.
@@
Consensus-update Actor-critic Algroithm

### v3

# Peer-to-Peer Energy Trading of Solar and Energy Storage: A Networked Multiagent Reinforcement Learning Approach

@@ Abstract
Utilizing distributed renewable energy resources, particularly solar and energy storage, in local distribution networks via peer-to-peer (P2P) energy trading has long been touted as a solution to improve energy systems’ resilience and sustainability. Consumers and prosumers (that is, those with solar PV and/or energy storage), however, do not have the expertise to engage in repeated P2P trading, and the zero-marginal costs of renewables present challenges in determining fair market prices. To address these issues, we propose multi-agent reinforcement learning (MARL) frameworks to help automate consumers’ bidding and management of their solar PV and energy storage resources, under a specific P2P clearing mechanism that utilizes the so-called supply-demand ratio. In addition, we show how the MARL frameworks can integrate physical network constraints to realize voltage control, hence ensuring physical feasibility of the P2P energy trading and paving way for real-world implementations.
keywords:
Multi-agent reinforcement learning, distributed energy resources, peer-to-peer market.
journal: Applied Energy
[…]
[label1]organization=School of Industrial Engineering, Purdue University,addressline=Grissom Hall, city=West Lafayette, postcode=47907, state=IN, country=USA
1 Introduction
@@
The rest of the paper is organized as follows. Section 2 reviews the literature on P2P energy trading and emphasizes the contribution of our work. Section 3 introduces the market clearing mechanism based on the supply-demand ratio, first proposed in [5], which can avoid zero-clearing prices when the supply resources have zero marginal costs. Section 4 provides the detailed consensus-MARL formulation and algorithm, as well as establishing convergence results. The specific setups and data of our simulation are described in Section 5, along with numerical results. Finally, Section 6 summarizes our work and points out several future research directions.
2 Literature Review
@@
The literature on decentralized control of energy transactions in a distribution network is extensive and diverse. Broadly speaking, there are two predominant approaches: the distributed optimization approach and P2P bilateral trading. A key method in the former approach is the alternating direction multiplier method (ADMM), which solves centralized dispatch problems by iteratively updating price signals for consumer/prosumer optimization (such as [6, 7, 8, 9, 10]). Despite ADMM’s theoretical convergence under convexity, practical challenges arise: (i) As an algorithmic approach rather than a market design, ADMM is subject to the same difficulties faced by a uniform-pricing-based market of all zero-marginal-cost resources. Specifically, it is prone to ‘bang-bang’ pricing: market clearing prices plummet to zero in oversupply or spike to a price ceiling in under-supply. Addressing this with additional market mechanisms, such as adding ancillary service markets, akin to wholesale markets, could overly complicate distribution-level markets for energy trading. (ii) Most distributed algorithms, including ADMM, are static, lacking real-time adaptability. While online ADMM variants exist [11, 12], ensuring network constraints within each iteration remains an open question. (iii) Assuming consumers or prosumers can solve complex optimization problems is often unrealistic due to potential expertise or resource limitations.
@@
These limitations form the basis of our interest in exploring P2P bilateral trading. Consequently, our literature review will focus on this alternative approach. There have been multiple review papers on the large and ever-growing literature on P2P energy trading [13, 14, 15, 16, 17]. Within the literature, we focus on the works that explicitly consider distribution network constraints. [18] proposes a continuous double auction framework within which physical network constraints are maintained through a sensitivity-based approach. The agents in the auction are the so-called zero-intelligence-plus (ZIP) traders, who employ simple adaptive mechanisms without any learning of the repeated multi-agent interactions. More sophisticated than ZIP, a fully decentralized MARL is adopted in [19] to realize decentralized voltage control and is extended in [2] to form a P2P energy trading market with physical constraints. These two papers implement a completely decentralized approach in the sense that each agent ignores the multi-agent interaction as well and just learns their single-agent policy to maximize their own payoff.
@@
Another MARL approach that has been explored is the centralized training and decentralized execution (CTDE) framework, of which the multi-agent deep deterministic policy gradient (MADDPG) method [4] is a prominent example. It is applied in [20] to solve the autonomous voltage control problem In [21], a modified MADDPG algorithm is proposed for a double-sided auction market to facilitate P2P energy trading, although physical constraints are not considered in their model. While the CTDE framework can lead to more efficient learning and robust outcomes, its dependence on a centralized entity for collecting and coordinating all agents’ data introduces difficulties for practical implementations and potential vulnerabilities, such as increased risk of single-point failures and higher susceptibility to cyber-attacks.
@@
In this work, we present several key contributions: First, we develop a comprehensive framework for a P2P energy trading market while integrating a consensus-based MARL algorithm. This framework facilitates automated control and bidding of DERs while allowing for decentralized learning of voltage constraints. Second, we present the sufficient conditions required for the convergence of the MARL algorithm and show the details of implementing this algorithm using deep neural networks. Third, we conduct a comparative analysis of market outcomes derived from three distinct MARL strategies: a naive decentralization approach, the MADDPG approach, and the proposed consensus-based MARL. Our findings reveal that the consensus-based approach notably outperforms the others in achieving better market outcomes. Our work addresses real-world applications of P2P energy trading under network constraints, which is crucial for integrating distributed renewable energy sources and enhancing grid stability. The RL-based algorithm we propose, designed for control automation, can be directly implemented on grid-edge devices, making it highly applicable for distribution networks with increasing numbers of prosumers.
3 Market Clearing Mechanism
@@
$P_{t}:=P(SDR_{t})=\begin{cases}(FIT-UR)\cdot SDR_{t}+UR,&0\leq SDR_{t}\leq 1;\\ FIT,&SDR_{t}>1.\end{cases}$ (2)
In (2), $P_{t}$ denotes the market clearing price in round $t$, which is a piece-wise linear function with respect to the $SDR_{t}$, as illustrated in Figure 1.
@@
While the SDR approach may be criticized for not being based on economic theories, we argue that its simplicity and transparency make it well-suited for a P2P energy market with zero marginal cost resources. In our previous work [24], we identified that a P2P energy market using a traditional double-sided auction is unlikely to perform well due to its unique characteristics. Specifically, uncleared electricity demand bids must be bought at the $UR$ rate, and uncleared energy supply bids must be sold at $FIT$, with both rates known to all market participants. The public information of $UR$ and $FIT$ naturally leads sellers to bid at $UR$ and buyers to bid at $FIT$, aiming to maximize their payoffs, since the marginal generation costs are zero. This will result in a bang-bang market outcome: when there is overdemand, the market price is $UR$; when there is oversupply, the price is $FIT$. We view the SDR approach as a fair and reasonable alternative to the marginal-cost-based approach to avoid such outcomes, especially in a new market where supply is likely much less than demand.
@@
Additionally, the SDR method offers two key benefits: First, it maintains the market clearing price between $UR$ and $FIT$, ensuring that transactions for buyers and sellers are at least as favorable as dealing directly with the utility. Second, it simplifies the bidding process by requiring only quantity bids. This allows consumers with flexible loads to focus on adjusting their demand timings, and prosumers to better manage their energy storage and sales strategies, without the complexity of submitting price-quantity pairs.
With the above being said, we are interested in comparing consumer and producer surplus (and their sum, social surplus) between the marginal-cost-pricing (MCP) approach and the SDR approach. Conceptually, consumer surplus is the difference between the maximum amount consumers are willing to pay for a good and the amount they actually pay. Similarly, producer surplus represents the difference between the price producers receive for a good and the cost of producing it. However, several complicating factors arise in surplus analysis here. First, both demand and supply have inflexible and flexible components. For electricity consumption, there is inflexible demand (the so-called baseload) and flexible demand, with flexibility either coming from shiftable loads (like EV charging, washers/dryers) or from charging energy storage. For baseload, we do not know the (aggregated) consumers’ true willingness to pay. We follow the approach in [25] and use the utility rate as a proxy for consumers’ willingness to pay when calculating consumer surplus for the inflexible load; that is, (UR - clearing price) $\times$ baseload. For flexible load, we assume consumers’ willingness to pay is represented by a downward-sloping curve, as shown in Figures 2 and 3. In these figures, $Q_{B}$ denotes the total baseload, and $\widehat{Q}$ denotes the total load, both flexible and inflexible. For supply, we have inflexible supply, such as grid-tied solar (that is, there is no storage). In this case, any energy generated beyond the prosumer’s own consumption is automatically sold back to the grid. We denote the total inflexible supply by $S_{B}$. There is also flexible supply, where flexibility comes entirely from the option to store energy for later sale. For flexible supply, we assume prosumers’ willingness to sell is represented by an upward-sloping curve until it hits the total supply capacity (denoted by $\widehat{S}$), as depicted in Figures 2 and 3.
Next, we analyze the differences in surpluses between MCP and SDR. We assume that under MCP, consumers and prosumers bid their true willingness-to-buy and willingness-to-sell curves, respectively. We must consider different cases we do not know how prosumers would bid under the SDR mechanism. The first case is when $\widehat{S}<Q_{B}$, meaning the maximum prosumers’ supply is less than the total baseload. In this scenario, under MCP, the market clearing price will be $UR$, and prosumers capture all the surplus (denoted as PS in the figures), as shown in the left graph of Figure 2. In the SDR case, regardless of the agents’ strategies, we will have $10<SDR<$ 1, since the total supply is less than demand. In this case, consumers capture the surplus of $(UR-P_{SDR})\times\widehat{S}$, as shown in the right graph of Figure 2, with the area denoted by CS, and it is easy to see that the social surplus is identical to the MCP case. This is likely the most common scenario now and in the near future, where prosumers’ total solar and storage capacity is less than total inflexible demand. The second case is the opposite extreme, where $S_{B}>\widehat{Q}$. In this case, $P_{SDR}=0$, which is the same as $P_{MCP}$. As a result, consumers capture all the surplus, and the social surplus is again identical between MCP and SDR, as shown in the left graph of Figure 3. In cases beyond these two extremes, the situation is less clear, as it depends on how agents bid in both the SDR and MCP markets. In such cases, we expect that $P_{SDR}$ and $P_{MCP}$ will differ, making the analysis of social surplus more complex. We defer this analysis to future research.
It is important to note, based on Figures 2 and 3, that the MCP approach would not present a problem if the willingness-to-buy and willingness-to-sell curves were known, even when all supply resources have zero marginal costs, as market clearing prices would not necessarily be 0 or $UR$. However, determining these curves is impractical, as even consumers and prosumers often do not know their precise willingness to buy or sell in relation to flexible demand and supply. This is exactly where the SDR approach offers a distinct advantage: it removes the need for agents to define their true preferences – which are inherently uncertain—and simply requires them to bid quantities. Moreover, in at least two cases, there is no loss in social surplus, and in fact, the SDR mechanism results in greater consumer surplus when demand significantly exceeds prosumer supply.
Figure 2: Social surplus comparison: total zero-marginal-cost supply less than total inflexible load
Figure 3: Social surplus: total zero-marginal-cost supply more than total inflexible load
4 Decentralized Learning with Consensus Updates
@@
4.1 Markov Decision Process for a Prosumer
@@
$e_{i,t+1}=\max\Big{\{}\min\Big{[}e_{i,t}+\eta^{c}_{i}\max(a^{e}_{i,t},0)+\frac{1}{\eta^{d}_{i}}\min(a^{e}_{i,t},0),\overline{e}_{i}\Big{]},0\Big{\}},$ (3)
where $\eta^{c}_{i}$ and $\eta^{d}_{i}$ are the charging and discharging efficiency of agent $i$’s battery, respectively, and $\overline{e}_{i}$ is the battery capacity. The charging efficiency represents the ratio of the amount of energy effectively stored in the system to the energy input during the charging process; while the discharging efficiency is the ratio of the energy delivered by the system during discharge to the energy that was initially stored in it. Their product, $\eta^{c}_{i}\times\eta^{d}_{i}$, yields the so-called round-trip efficiency. The outside “max” operation in (3) is to ensure that the energy level in the battery will not be negative; while the first “min” operation in the bracket is to ensure that the battery storage capacity is not exceeded.
Reward: Each agent $i$’s reward function is affected by both the collective actions of the agents and the system states. Specifically, agent $i$’s reward in time step $t$, denoted by $r_{i,t}$, is a random variable whose conditional expectation has three components as follows:
$\begin{split}\mathbb{E}[r_{i,t}|s_{t},a_{t}]:=&\ \mathbb{E}[r^{m}_{i}(a^{e}_{i,t};a^{e}_{-i,t},s_{t})]+\mathbb{E}[r^{v}(a_{i,t};a_{-i,t},s_{t})]/I\\[5.0pt] &+\mathbb{E}[r^{l}(a_{i,t};a_{-i,t},s_{t})]/I.\end{split}$ (4)
@@
which is simply the net energy of PV generation minus baseload demand (of real power) and charge to the battery. Since PV generation is inherently variable, this net available energy is also a random variable. The market clearing price $P_{t}$ in (5) is defined in (2).
@@
The second term in the reward function (4), $r^{v}$, is the system voltage violation penalty at a given time $t$. To provide an explicit form of $r^{v}$, we first write out the standard bus-injection model. Assume that the network has $N$ buses, we have that for $\kappa=1,\ldots,N$ (the time index $t$ is omitted below for simplicity):
@@
In (8), $\overline{V}_{\kappa}$ and $\underline{V}_{\kappa}$ represent the upper and lower limit of bus $k$’s voltage magnitude, respectively. Let $f_{clip}$ denote a generic ‘clip’ function in the form of $f_{clip}=\text{clip}[x,\{v_{min},v_{max}\}]$, which is defined as $f_{clip}=x$ when $v_{min}\leq x\leq v_{max}$, $f_{clip}=v_{min}$ if $x<v_{min}$, and $f_{clip}=v_{max}$ if $x>v_{max}$. Based on the clip function definition, if the voltage magnitudes $|V_{\kappa,t}|$ obtained from solving equations (7) are within the voltage limit of $[\underline{V}_{\kappa},\overline{V}_{\kappa}]$ on every bus $\kappa=1,\ldots,N$, then $R^{v}=0$; if there is voltage violation at a certain bus $\kappa$, the term $\max(|V_{\kappa,t}|-\overline{V}_{\kappa},\ \underline{V}_{\kappa}-|V_{\kappa,t}|)$ becomes positive, and $r^{v}$ becomes a negative number with an arbitrary positive number $\lambda$ to amplify the voltage violation. The other positive parameter in (8), $M$, is just to ensure that the reward $r^{v}$ is bounded below, a technical condition needed to ensure the convergence of the consensus MARL algorithm to be introduced in the next section. If the power flow equations (7) do not have a solution with a given set of agents’ bids, we can simply set $r^{v}=-\lambda\cdot N\cdot M$.
The third term in the reward function (4), $r^{l}$, is the line capacity violation penalty over the entire distribution network at a given time $t$. Let $S_{\kappa,j}$ denote the apparent power flow from bus $\kappa$ to bus $j$, which is defined as $S_{\kappa,j}=V_{j}I_{\kappa,j}^{*}$, where $I_{\kappa,j}^{*}$ is the complex conjugate of $I_{\kappa,j}$ – the current from bus $\kappa$ to bus $j$. $I_{\kappa,j}$ is calculated $I_{\kappa,j}=Y_{\kappa,j}(V_{\kappa}-V_{j})$, where $Y_{\kappa,j}$ is the complex admittance of the branch between buses $\kappa$ and $j$, and $V_{\kappa}$ and $V_{j}$ are the voltages from the power flow equations in (7). Let $\mathcal{E}$ be the set of branches with constraints on the magnitude of the apparent power flow. The total line capacity violation penalty can then be defined as follows:
$r^{l}:=-\gamma\sum_{(\kappa,j)\in\mathcal{E}}\min\Big{[}\max(|S_{\kappa,j}|-\overline{S}_{\kappa,j},0),C\Big{]},\vspace*{-5pt}$ (9)
where $\overline{S}_{\kappa,j}$ represent the upper limit of branch $\kappa-j$, $C$ is a constant to ensure that the reward $r^{l}$ is bounded below, and $\gamma$ is an arbitrary positive number to amplify the capacity violation. If (7) does not have a solution, similar to the voltage violation case, we set $r^{l}=-\lambda\cdot N\cdot C$.
We want to clarify that the voltage and line capacity violation penalties are not actual financial penalties for the agents, as market clearing occurs before real-time energy delivery. Instead, these penalties only serve as feedback to help agents refine their reinforcement learning policies. Note that the (fictitious) network violation penalties should not be confused with the fees utilities charge for grid maintenance. This trading market does not account for how utilities impose transmission and distribution (T&D) maintenance fees, which are typically fixed rather than volumetric in most U.S. regions [29]. We can adjust the utility rate (UR, the price ceiling in Fig. 1) to subtract the fixed charge, and the framework in this paper remains applicable, with the understanding that each transaction will have to pay an additional grid maintenance fee.
@@
4.2 Consensus-update Actor-critic Algroithm